OpenStudy (ehall):

If the sin A =3/5 and the cos A =4/5 , then what is tan A? 3/4 1 4/3

OpenStudy (holsteremission):

\[\tan A=\frac{\sin A}{\cos A}=\cdots\]

OpenStudy (ehall):

i got what do i do next

OpenStudy (ehall):

please help me

OpenStudy (karim728):